04 July 2014 UT Mini-queue night

Summary

Night delayed due to heavy, threatening clouds. Did not open for OSURC portion of night

Plan

Hoped to start with LBC on high-priority UVa program, but thick clouds delayed and then prevented opening. Turnover to Pathfinder around 12:30.

Log

02:30 - Clouds are very thick. Opening is delayed.

03:30 - Thicker clouds settling in. Satellite is terrible.

04:30 - Still really cloudy!

05:30 - Ditto

07:00 - Giving up the OSURC queue for the night. Turning over to Geno and then Pathfinder at 12:30.

Pathfinder log

07:30 - raining

08:00 - no rain

08:40 - can see summer triangle

09:00 - clouds clearing up

09:45 - weather would allow to open the dome

09:45 - some technical issues with 208 drive cabinet - master power reset (?) pressed

09:45 - reset apparently caused technical problems with the computer of the Telescope Operator. Reboot required!

10:05 - setting up the operator workspace with remote assistance by Doug Fischer.

10:20 - apparently also other systems are affected. OBS1. TCS control GUI is all blue. Hexapods are red. Error message continuously displayed which cannot be interpreted by Geno nor by Doug.

10:25 - Geno is still trying to get the telescope operational

10:40 - Remote assistance by Michelle

11:05 - Michelle diagnosed the problem. Network failure. Not all systems were properly shut down. Michelle is manually shutting down each system.

11:06 - it is getting more and more cloudy.

11:36 - Telescope is still not operational. Giving up for tonight.
Topic revision: r4 - 06 Jul 2014, ThomasBertram
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback